OpenStudy (anonymous):

How do I find the EXACT value of the following. 1. sin (pi/12) cos (5 pi/12) + cos (pi/12) sin (5pi/12) 2. cos (2pi/3 - pi/6) 3. sin (5pi/12) For number two I know you have to do something with identiden

OpenStudy (anonymous):

1. \(\sin (\pi/12 + 5\pi/12)\) using: sin (a+b)

OpenStudy (anonymous):

\[\sin \left( A+B \right)=\sin A \cos B+\cos A \sin B\]\[\cos \left( A-B \right)=\cos A \cos B+\sin A \sin B\]

OpenStudy (anonymous):

Ah yes I just got that one. It's 1!

OpenStudy (anonymous):

Actually it's 1.2247

OpenStudy (anonymous):

For number three is it 3.647?

OpenStudy (anonymous):

I just got number 4.